OpenStudy (anonymous):

area of triangle the base is 2 1/3 yrds and the height is 3/4

OpenStudy (anonymous):

please help me

OpenStudy (raden):

use formula : Area (A) = b*h/2

OpenStudy (anonymous):

half of base multiplied by height

OpenStudy (anonymous):

but it is a fraction so how would I figure out 1/2 the base

OpenStudy (anonymous):

2 1/3 is 7/3

OpenStudy (anonymous):

7/3*3/4 right

OpenStudy (anonymous):

1/2 * 7/3 = 17/6 look up mixed numbers and multiplying fractions google this patrik jmt

OpenStudy (anonymous):

you need to find the lowest common denominator which is 6

OpenStudy (anonymous):

then what you do to the bottom you have to do to the top

OpenStudy (anonymous):

so 1/2 * 7/3 1/6 * 7/6 but you have to multiply the top as well so it becomes 3/6 *14/6

OpenStudy (anonymous):

42/6=7

OpenStudy (anonymous):

so 17/6 * 3/4 same process 17/12 * 3/12 but you have to multiply the top as well 51/12 * 9/12 = 60/12 this should be the answer

OpenStudy (anonymous):

60/12=5 yd^2

OpenStudy (anonymous):

it says for the answer ?? yrds ^2

OpenStudy (anonymous):

g.o.a.t

OpenStudy (anonymous):

muntoo I am going in sane

OpenStudy (anonymous):

Remember the formula: area = base x height 2 1/3 yrds is the same thing as 7/3 yrds. area = 7/3 x 3/4

OpenStudy (anonymous):

i do not know how to do this goat confused me more

OpenStudy (anonymous):

so the bottom is 12 right

OpenStudy (anonymous):

For your other questions, remember to write down every single step. Don't skip steps yet. Go over what you write to make sure you haven't made any mistakes.

OpenStudy (anonymous):

Sorry, I messed up: area = base x height / 2 2 1/3 yrds is the same thing as 7/3 yrds. area = (7/3) x (3/4) / 2

OpenStudy (anonymous):

That's the formula for triangles only.

OpenStudy (anonymous):

I know but how do I figure that out on calculator

OpenStudy (anonymous):

If you're using a "simple" calculator: 7 divide 3 multiply 3 divide 4 divide 2

OpenStudy (anonymous):

.875

OpenStudy (anonymous):

I feel so stupid, I have been doing 50 problems for 11 hours and still have some to go. I have to get this our I will fail my aa degree

OpenStudy (anonymous):

am I right on the answer

OpenStudy (anonymous):

on the answer it is saying yrd^2

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

sorry 17/6 should of been 7/6

OpenStudy (anonymous):

carried the same mistake all the way through on the multiplication of fractions

OpenStudy (anonymous):

so my answer is .875 yrds^2

OpenStudy (anonymous):

area = (7/3) yrd x (3/4) yrd / 2 = 0.875 yrd yrd = 0.875 yrd^2 Yes.

OpenStudy (anonymous):

should end up with 14/12 * 9/12 = 7/8 sorry bout the errors

OpenStudy (anonymous):

i will start again so u have full workings to help.

OpenStudy (anonymous):

thanks muntoo and goat

OpenStudy (anonymous):

I have never been so workup on math like this before, I am bolling my eyes out bevause I feel like I knew it when I started this class and now I can not remember nothing

OpenStudy (anonymous):

Tips: 1) Show all your work. Be organized. 2) Study a little every day. A few days later (i.e. two days) go over what you learned that day and see if you remember. At the end of the week do a review of what you learned that week, and don't learn anything new that day. The key is not to do everything in one day -- just work on it every once in a while till you get it.

OpenStudy (anonymous):

this is the final and I only get one chance that is why I am still on here because I can not get it but I cant stop because of my one chance